# Symmetry verification for noisy quantum simulations of non-Abelian lattice gauge theories

### v1

@@ Appendix C $D_{3}$ lattice gauge theory
Figure 8: Comparison between the exact, noisy, DPS, and PSV time evolution in presence of dephasing noise for (a) two plaquettes, averaged over 5000 independent trajectories, and (b) three plaquettes, averaged over 145 independent trajectories. (c) At constant local error rate, the weight of the gauge-invariant part of the state for PSV and the survival probability in the case of DPS remain comparable (for DPS, there is still a single measurement per Trotter step). Thus, from 2 to 3 plaquettes the protection quality of DPS only slightly decreases, due to the reduced measurement density, while maintaining a constant resource cost. For PSV, by contrast, the protection quality even increases, at the cost of 6 times more observables to be measured, and thanks to the smaller ratio between the dimensions of the physical sector and the total Hilbert space.
@@ Appendix D Error model
Several techniques can numerically generate random unitaries close to the identity. Here, we implement a method based on Householder transformations [76]. These are unitary transformations describing the reflection about a hyperplane that contains the origin. The matrix operator describing the Householder reflection defined by its normal vector $\bm{v}$ reads $\mathcal{R}=\mathbb{1}-2\bm{v}\bm{v}^{\dagger}$, where $\bm{v}$ in our case is a random complex vector whose elements are extracted from a normal distribution with unit variance.
@@ Appendix E Gauge protection on three plaquettes
The overall picture continues to hold when considering the larger system with 3 plaquettes, as reported in Fig. 8(b), where we again monitor the left plaquette of the lattice. In DPS, we again measure only one local symmetry after each Trotter step so that the overall cost of the mitigation technique does not increase with respect to the two-plaquette model. Remarkably, the reduced measurement density only slightly affects the gauge protection, compared to the smaller system shown in Fig. 8(a). For a fair comparison, the noise strength here is rescaled such that the overall error rate $\sim\gamma^{2}$ is proportional to the system size. Given a small local error probability $p_{\rm err}$, the error rate per Trotter step for $N$ qudits is approximately given by $\gamma^{2}\simeq 1-(1-p_{\rm err})^{N}\simeq Np_{\rm err}$. Hence, when going from two to three plaquettes, we change the global error rate as $\gamma^{2}_{3\mathrm{P}}=3\gamma^{2}_{2{\rm P}}/2$, to keep a constant local error rate at first order in $p_{\rm err}$. Because the measurement density is also rescaled down with the same ratio $3/2$, the decay of the DPS survival probability behaves similarly in the two cases, as shown in the right panel of Fig. 8(c).
Notably, PSV now better approximates the exact time evolution with respect to the data shown for the two-plaquette lattice. This counterintuitive behavior can be explained by considering the ratio between the dimensions of the physical sector and the total Hilbert space. In the two-plaquette lattice, one has $\frac{\textrm{dim}\mathcal{H}_{\textrm{P}}}{\textrm{dim}\mathcal{H}_{\textrm{tot}}}=\frac{49}{1296}\simeq 0.0378$, while in the three-plaquette system $\frac{\textrm{dim}\mathcal{H}_{\textrm{P}}}{\textrm{dim}\mathcal{H}_{\textrm{tot}}}=\frac{251}{46656}\simeq 0.0054$. The smaller fraction reduces the probability of effective gauge-invariant errors originating from two consecutive gauge-violating errors that cancel, which PSV is not able to reveal. Hence, the fraction of noisy states projected into the physical sector retains information about the physical dynamics for a longer time. However, considering the overhead scaling with the number of vertices discussed in Sec. IV.2.2, suggests DPS could hold a practical advantage if mid-circuit measurements are available with sufficient speed and accuracy. Interestingly, the weight of the symmetrized density matrix $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]$ also does not display appreciable differences between the two- and three-plaquette ladders, see left panel of Fig. 8(c).

### v2

@@ Appendix C $D_{3}$ lattice gauge theory
DPS and PSV with qubits
Even though in this work we explore the application of DPS and PSV on a qudit-based hardware, these methods can be used on a qubit-based platform as well. In DPS, employing qubit-based hardware would increase the number of entangling gates to perform the change of basis in Fig. 3, an operation that on qudits can be implemented with single-qudit rotations. The resources needed depend on the group structure and on the native gates of the qubit hardware. Furthermore, the number of entangling gates for the encoding operation in Fig. 7 increases as well. For example, for $D_{3}$, the verification of the rotational symmetries will require entangling two auxiliary qubits with the $12$ qubits used to encode the degrees of freedom around a vertex. For the PSV, using qubits instead of qudits requires an increasing number of entangling gates for the change of basis before the measurement.
@@
Figure 8: Comparison between the exact, noisy, DPS, and PSV time evolution in presence of dephasing noise for (a) two plaquettes, averaged over 5000 independent trajectories, and (b) three plaquettes, averaged over 145 independent trajectories. (c) At constant local error rate, the weight of the gauge-invariant part of the state for PSV and the survival probability in the case of DPS remain comparable (for DPS, there is still a single measurement per Trotter step). Thus, from 2 to 3 plaquettes the protection quality of DPS only slightly decreases, due to the reduced measurement density, while maintaining a constant resource cost. For PSV, by contrast, the protection quality even increases, at the cost of 6 times more observables to be measured, and thanks to the smaller ratio between the dimensions of the physical sector and the total Hilbert space. For the two-plaquette system, $\gamma=0.2$ as previously, while for the three-plaquette system we rescaled to the corresponding $\gamma=0.245$, as explained in the text.
@@ Appendix D Error model
Several techniques can numerically generate random unitaries close to the identity. Here, we implement a method based on Householder transformations [78]. These are unitary transformations describing the reflection about a hyperplane that contains the origin. The matrix operator describing the Householder reflection defined by its normal vector $\boldsymbol{v}$ reads $\mathcal{R}=\mathbb{1}-2\boldsymbol{v}\boldsymbol{v}^{\dagger}$, where $\boldsymbol{v}$ in our case is a random complex vector whose elements are extracted from a normal distribution with unit variance.
@@
A realistic estimate for the error rate strongly depends on the individual qudit hardware. Reference [59] also consider a small lattice with a pure gauge theory and a discrete group, where the main sources of error are control-multiplications used to implement the plaquette operators. On a Rydberg-atom qudit simulator, they estimate the achievable fidelity to be $99.6\%$. Since six of such two-qudit gates are required for each plaquette operator, one can expect the overall fidelity of a single Trotter step to be about $95\%$ for the two-plaquette lattice we have considered in this work. This corresponds to a noise strength $\gamma\sim 0.22$, which is compatible with the regime in which both PPS and DPS are found to work. At the same time, the bare noisy dynamics quickly relaxes to a nonphysical steady state, highlighting the necessity of error-mitigation schemes.
@@ Appendix E Gauge protection on three plaquettes
The overall picture continues to hold when considering the larger system with 3 plaquettes, as reported in Fig. 8(b), where we again monitor the left plaquette of the lattice. In DPS, we again measure only one local symmetry after each Trotter step so that the overall cost of the mitigation technique does not increase with respect to the two-plaquette model. Remarkably, the reduced measurement density only slightly affects the gauge protection, compared to the smaller system shown in Fig. 8(a). For a fair comparison, the noise strength here is rescaled such that the overall error rate $\sim\gamma^{2}$ is proportional to the system size. Given a small local error probability $p_{\rm err}$, the error rate per Trotter step for $N$ qudits at first order in $p_{\rm err}$ is given by $\gamma^{2}\simeq 1-(1-p_{\rm err})^{N}\simeq Np_{\rm err}$. Here, $p_{\rm err}$ reflects the average fault rate per qudit resulting from all gates applied in the implementation of a single Trotter step. When increasing the system size from two to three plaquettes, we assume that this microscopic error rate $p_{\rm err}$ remains fixed. With the number of operations per layer growing linearly with the number of plaquettes, this leads to a corresponding linear increase of the overall infidelity $\gamma^{2}$ for each Trotter step, i.e. $\gamma^{2}_{3\mathrm{P}}=3\gamma^{2}_{2{\rm P}}/2$. Because the measurement density is also rescaled down with the same ratio $3/2$, the decay of the DPS survival probability behaves similarly in the two cases, as shown in the right panel of Fig. 8(c).
Notably, PSV now better approximates the exact time evolution with respect to the data shown for the two-plaquette lattice. This counterintuitive behaviour can be explained by considering the ratio between the dimensions of the physical sector and the total Hilbert space. In the two-plaquette lattice, one has $\frac{\textrm{dim}\mathcal{H}_{\textrm{P}}}{\textrm{dim}\mathcal{H}_{\textrm{tot}}}=\frac{49}{1296}\simeq 0.0378$, while in the three-plaquette system $\frac{\textrm{dim}\mathcal{H}_{\textrm{P}}}{\textrm{dim}\mathcal{H}_{\textrm{tot}}}=\frac{251}{46656}\simeq 0.0054$. The smaller fraction reduces the probability of effective gauge-invariant errors originating from two consecutive gauge-violating errors that cancel, which PSV is not able to reveal. Hence, the fraction of noisy states projected into the physical sector retains information about the physical dynamics for a longer time. However, considering the overhead scaling with the number of vertices discussed in Sec. 4.2.2 suggests DPS could hold a practical advantage if mid-circuit measurements are available with sufficient speed and accuracy. Interestingly, the weight of the symmetrized density matrix $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]$ also does not display appreciable differences between the two- and three-plaquette ladders, see left panel of Fig. 8(c).